Solve the following system of equations:
9x+5y=-4
-8x+3y=7

Found 2 solutions by kekegeter , Alan3354:
Answer by kekegeter (97)

9x+5y=-4
9x=-5y-4
x=-5y+4/9

8x+3y=7
-8x-3y+7
x=- -3y+7/8

Answer by Alan3354(69443)

Solve the following system of equations:
9x+5y=-4
-8x+3y=7
----------------
Get the coefficients of x or y to match.
----
27x + 15y = -12 Eqn1 times 3
-40x + 15y = 35 Eqn2 times 5
----------------------------------- Subtract
67x = -47
x = -47/67
------------
Sub for x in any eqn to find y.
-8x+3y=7
376/67 + 3y = 7
376/67 + 3y = 469/67
3y = 93/67
y = 31/67